1) the sum of the first 12 of an A.p is -96 and the Sum of the first 18th term is -252. find the
(a)common difference and
(b)the first term
2) How many terms of an A.p, 5,13,21............ give as sum of 275

Answer by Boreal(15235)

sum is (n/2)(a1+an)
for 12 terms 6(a1+a12)=-96
for 18 terms 9*(a1+a18)=-252
a1+a12=-16
a1+a18=-28
d=-2 ANSWER
sum = (n/2)(2a1+(n-1)d)
=6(2a1+11*-2)=-96
-16=2a1-22
2a1=6
a1=3 ANSWER
sum = -96=(12/2)(a1+a12)
-16=3+a12
a12=-19
a18=-31
==================
there is not an exact number of terms
5,13,21,29,37,45,53,61,69
sum=(n/2)(a1+an)
when n=8 and an=a1+(n-1)8, with difference of 8
an=5+56=61
and sum is 4(61) or 244, and the next term isn't 31